As most of you probably know the number of hours of sunlight in a day fluctuates thought the year. Each year has one shortest day and one longest day. This year in Eugene Oregon the longest day will be June 20 and the shortest day of the year is December 20. For simplicity assume that the shortest day has 6 hours of day light and the longest day has 18 hours of day light. Use the general sine or cosine function to module the number of hours of daylight as a function of the day of the year.
Answer #1

h = Asin (Bd + C + D h is wo. of hours of daylight d = day hmasi = D + A = 18 hmin. = D-A = 6 (d=01, 1st January) Hence, D =